OpenStudy (coolguy34):

I need help solving this equation: -2/5x - 9 < 9/10 please show me the steps

OpenStudy (coolguy34):

please i have been working on this all day

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[-\frac{2}{5}x-9<\frac{9}{10}\]?

OpenStudy (coolguy34):

yes

OpenStudy (misty1212):

ok lets get rid of the stupid fractions b multiplying EVERYTHING by 10 to start with \[-4x-90<9\]

OpenStudy (coolguy34):

ok

OpenStudy (misty1212):

add 90 to both sides \[-4x<99\]

OpenStudy (misty1212):

then divide by 4 and change the inequality from less than to greater than

OpenStudy (misty1212):

oops i meant "divide by \(-4\)"

OpenStudy (misty1212):

\[x>-\frac{99}{4}\]

OpenStudy (coolguy34):

ok

OpenStudy (coolguy34):

thanks